# PEA-Diffusion: Parameter-Efficient Adapter with Knowledge Distillation in non-English Text-to-Image Generation

### v1

@@ 5 Conclusion and Limitation
Conclusion. The expensive training costs and scarce training data for multilingual T2I models have led us to seek a lightweight solution. In this paper, we propose a simple and efficient approach through the introduction of KD and adapters. By only training the adapter while keeping other parameters frozen, our approach taps into the potential multilingual cultural understanding and generation capabilities of T2I models. Our approach achieves significantly better performance for culturally relevant data and outperforms existing models. It also maintains general synthesis ability. We conduct extensive experiments, demonstrating that our approach takes into account low-cost, versatility, and plug-and-play capabilities.
Limitation. Firstly, our approach requires replacing the CLIP text encoder with the corresponding language. The performance of the CLIP text encoder largely determines the quality of image generation and the relevance of the text. The weaker language-specific CLIP will affect PEA-Diffusion’s performance in general open domains and may result in a small sacrifice of the model’s generalization capabilities. Secondly, the ceiling of our approach depends on the English base model’s capacity, meaning that our approach cannot surpass its general capabilities. In the future, we will delve into further research to address these limitations.
Supplementary Material
Firstly, in Appendix 6, we provide information on how we organize and prepare the training data. We also present URL links to all publicly available models and code used in the paper.
Next, in Appendix 7, we analyze the performance of the Chinese language model with different training data resolutions and present detailed quantitative analysis of all languages in both the general evaluation set (MG) and the culture-specific evaluation set (MC) as the number of training samples increases. Finally, we show qualitative analysis in other languages.
In Appendix 8, we provide the results of the ablation experiments on hybrid training and details of ablation experiments in Sec. 4.4.
Appendix 9 analyzes low-cost experiments.
Finally, in Appendix 10, we provide additional visualizations of down-stream text-to-image generation tasks in Chinese.
@@ 6.2 Model or Code Path
The datasets, models and codes used in this paper are listed below, including Laion2B-multi [^1] https://huggingface.co/datasets/laion/laion2B-multi-joined-translated-to-en, Chinese CLIP [^2] https://github.com/OFA-Sys/Chinese-CLIP, Russian CLIP [^3] https://github.com/ai-forever/ru-clip, Japanese CLIP [^4] https://github.com/rinnakk/japanese-clip, Korean CLIP [^5] https://github.com/jaketae/koclip, and Italian CLIP [^6] https://github.com/clip-italian/clip-italian, ProtoVision XL [^7] https://civitai.com/models/125703/protovision-xl-high-fidelity-3d-photorealism-anime-hyperrealism-no-refiner-needed, OpenCLIP ViT-bigG [^8] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k, SD1.5999https://huggingface.co/runwayml/stable-diffusion-v1-5 , AltDiffusion-m9101010https://huggingface.co/BAAI/AltDiffusion-m9, AltDiffusion-18111111https://huggingface.co/BAAI/AltDiffusion-m18, GlueGen code [^12] https://github.com/salesforce/GlueGen, “xxmix9realistic” [^13] https://civitai.com/models/124421/xxmix9realisticsdxl, “samaritan-3d-cartoon” [^14] https://civitai.com/models/81270/samaritan-3d-cartoon, “csal_scenery” [^15] https://civitai.com/models/118559/ancient-chinese-scenery-background-xl, “controlnet-canny” [^16] https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, SD-XL Inpainting 0.1171717https://huggingface.co/diffusers/stable-diffusion-xl-1.0-inpainting-0.1, Latent Consistency Models[20] [^18] https://huggingface.co/latent-consistency/lcm-lora-sdxl, Segmind Stable Diffusion Model [^19] https://huggingface.co/segmind/SSD-1B. In addition, we also plan to release our code.
@@ 7.2.1 Quantitative Analysis with Other Language
Table 3 shows that Russian is the second largest language data in our collection. On the general evaluation dataset MG. Figure 5 shows the ImageReward score for Russian data can remain stable at around 0.77, which is higher than the results of training Chinese data. Moreover, Table 1 indicates that the CLIPScore on the MC evaluation dataset for Russian is higher than the results with the translation-based model by approximately 2.5 percentage points.
@@
Table 3 shows that the amount of Korean data used for training is only 1.6M. Figure 6 shows the ImageReward score on the general evaluation dataset MG can still remain stable at around 0.43. Additionally, Table 1 shows that on the MC evaluation dataset for Korean, the CLIPScore is higher than the results with the translation-based model by approximately 1.5 percentage points.
@@
Table 3 shows that the amount of Italian data used for training is only 18M. Figure 7 shows the ImageReward score on the general evaluation dataset MG can still remain stable at around 0.3. Additionally, Table 1 shows that on the MC evaluation dataset for Italian, the CLIPScore is higher than the results with the translation-based model by approximately 5 percentage points.
@@ 7.2.2 Quantitative Analysis with Specified CLIPScore
As shown in Table 4, we also calculate the CLIP scores of the five languages using the CLIP model with corresponding languages in the MG teseset. It can be observed that out PEA-Diffusion still outperforms translation methods in terms of Chinese, Italian and Japanese while is only slightly inferior on Russian and Korean. This phenomenon demonstrates that translation method will lose critical information even for general prompts, leading to image-text matching degradation. Our PEA-Diffusion avoids the translation procedure while generates images based on the original prompts by an end-to-end manner, which will be a more optimal strategy.
@@ 9.1 Analysis of PEA Parameter
9.2 Fine-tuning
As a low-cost experiment for comparison, we also finetune the model to transfer language capabilities. Finetuning an image generation model can often lead to catastrophic forgetting, which significantly degrade the model’s open-domain performance. To address these issues, we adopt a parameter-efficient finetuning strategy that preserves the original model’s ability to the greatest extent possible. During the learning stage, we only update the network parameters required for learning language character generation while freezing other parameters. We adopt two fine-tuning schemes here. The first on is Fine-tuning K and V matrices of the attention laters in the UNet. As proposed in [15], when updating the mapping from given text to image distribution, only updating $W_{K},W_{V}$ in each cross-attention block $i$ is sufficient since text features are the only input to the key and value projection matrix. The second one is fine-tuning with LoRA [12].
Evaluation only $W_{k}^{(i)}$ and $W_{v}^{(i)}$ LoRA PEA-Diffusion
CLIPScore_zh 0.2485 0.2366 0.2610
ImageReward 0.0293 -0.3214 1.0245
CLIPScore_en 0.3854 0.3577 0.4224
PickScore 0.1758 0.1613 0.2244
@@ 9.2 Fine-tuning
10 Downstream Experiment
10.1 LCM Quantitative Analysis
Latent Consistency Models [19] have achieved impressive performance in accelerating T2I generative tasks, producing high-quality images with minimal inference steps. “lcm-lora-sdxl” is a universal acceleration module based on SDXL that allows to reduce the number of inference steps to only between 2 and 8 steps.
10.2 More Downstream Generated Images
Figure 13 shows more image generation situations of downstream adaptation. The examples given here are still in Chinese.

### v2

@@ 5 Conclusion and Limitation
Conclusion. The high training expenses and limited training data available for multilingual T2I models have driven us to search for a more lightweight solution. In this paper, we introduce a straightforward and efficient method using KD and adapters. By focusing solely on training the adapter and keeping other parameters fixed, our approach harnesses the capability of T2I models for multilingual cultural comprehension and generation. Our method not only achieves significantly better results for culturally relevant data but also surpasses existing models while maintaining general synthesis capability. Comprehensive experiments demonstrate that our method is cost-effective, versatile, and easy to integrate.
Limitation. First, the effectiveness of the CLIP text encoder significantly influences both the quality of image generation and text relevance. Utilizing a weaker language-specific CLIP could impact the performance of PEA-Diffusion in broad open domains and may slightly compromise the model’s generalization abilities. Second, the upper limit of our method is constrained by the capabilities of the English base model, indicating that our approach cannot exceed its inherent potential. Moving forward, we aim to conduct additional research to mitigate these limitations.
Appendix
Appendix 6 provide information on how we organize and prepare the training data. We also present URL links to all publicly available models and code used in the paper.
Appendix 7 analyze the performance of the Chinese language model with different training data resolutions and present detailed quantitative analysis of all languages in both the general evaluation set (MG) and the culture-specific evaluation set (MC) as the number of training samples increases. Finally, we show qualitative analysis in other languages.
Appendix 8 provide the results of the ablation experiments on hybrid training and details of ablation experiments in Sec. 4.4.
Appendix 9 analyze low-cost experiments.
Appendix 10 analyze the performance of LLMs as text encoder for T2I.
Appendix 11 analyze the qualitative results for Sec. 4.3 Table 1 in the general domain.
Appendix 12 analyze the user study.
Appendix 13 provide additional visualizations of down-stream text-to-image generation tasks in Chinese.
@@ 6.2 Model or Code Path
The datasets, models and codes used in this paper are listed below, including Laion2B-multi [^0] https://huggingface.co/datasets/laion/laion2B-multi-joined-translated-to-en, Chinese CLIP [^1] https://github.com/OFA-Sys/Chinese-CLIP, Russian CLIP [^2] https://github.com/ai-forever/ru-clip, Japanese CLIP [^3] https://github.com/rinnakk/japanese-clip, Korean CLIP [^4] https://github.com/jaketae/koclip, and Italian CLIP [^5] https://github.com/clip-italian/clip-italian, ProtoVision XL [^6] https://civitai.com/models/125703/protovision-xl-high-fidelity-3d-photorealism-anime-hyperrealism-no-refiner-needed, OpenCLIP ViT-bigG [^7] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k, SD1.5888https://huggingface.co/runwayml/stable-diffusion-v1-5 , AltDiffusion-m9999https://huggingface.co/BAAI/AltDiffusion-m9, AltDiffusion-18101010https://huggingface.co/BAAI/AltDiffusion-m18, GlueGen code [^11] https://github.com/salesforce/GlueGen, “xxmix9realistic” [^12] https://civitai.com/models/124421/xxmix9realisticsdxl,
“samaritan-3d-cartoon” [^13] https://civitai.com/models/81270/samaritan-3d-cartoon, “csal_scenery” [^14] https://civitai.com/models/118559/ancient-chinese-scenery-background-xl, “controlnet-canny” [^15] https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, SD-XL Inpainting 0.1161616https://huggingface.co/diffusers/stable-diffusion-xl-1.0-inpainting-0.1, Latent Consistency Models[21] [^17] https://huggingface.co/latent-consistency/lcm-lora-sdxl, Segmind Stable Diffusion Model [^18] https://huggingface.co/segmind/SSD-1B. SDXL-Turbo Model [^19] https://huggingface.co/stabilityai/sdxl-turbo. ChatGLM-6B  [^20] https://huggingface.co/THUDM/chatglm3-6b, Qwen-7B  [^21] https://huggingface.co/Qwen/Qwen-7B, Baichuan2-7B  [^22] https://huggingface.co/baichuan-inc/Baichuan2-7B-Base, Palyground V2  [^23] https://huggingface.co/playgroundai/playground-v2-1024px-aesthetic, Kandinsky V3  [^24] https://huggingface.co/kandinsky-community/kandinsky-3, Stable Cascade  [^25] https://huggingface.co/stabilityai/stable-cascade, In addition, we also plan to release our code.
@@ 7.2.1 Quantitative Analysis with Other Language
Table 4 shows that Russian is the second largest language data in our collection. On the general evaluation dataset MG. Figure 9 shows the ImageReward score for Russian data can remain stable at around 0.77, which is higher than the results of training Chinese data. Moreover, Sec. 4.3 Table 1 indicates that the CLIPScore on the MC evaluation dataset for Russian is higher than the results with the translation-based model by approximately 2.5 percentage points.
@@
Figure 9: Japanese evaluation results.
Table 4 shows that the amount of Korean data used for training is only 1.6M. Figure 9 shows the ImageReward score on the general evaluation dataset MG can still remain stable at around 0.43. Additionally, Sec. 4.3 Table 1 shows that on the MC evaluation dataset for Korean, the CLIPScore is higher than the results with the translation-based model by approximately 1.5 percentage points.
Table 4 shows that the amount of Italian data used for training is only 18M. Figure 9 shows the ImageReward score on the general evaluation dataset MG can still remain stable at around 0.3. Additionally, Sec. 4.3 Table 1 shows that on the MC evaluation dataset for Italian, the CLIPScore is higher than the results with the translation-based model by approximately 5 percentage points.
@@ 7.2.2 Quantitative Analysis with Specified CLIPScore
As shown in Table LABEL:tab:CLIP_five, we also calculate the CLIP scores of the five languages using the CLIP model with corresponding languages in the MG teseset. It can be observed that out PEA-Diffusion still outperforms translation methods in terms of Chinese, Italian and Japanese while is only slightly inferior on Russian and Korean. This phenomenon demonstrates that translation method will lose critical information even for general prompts, leading to image-text matching degradation. Our PEA-Diffusion avoids the translation procedure while generates images based on the original prompts by an end-to-end manner, which will be a more optimal strategy.
@@ 9.1 Analysis of PEA Parameter
9.2 Fine-tuning
As a low-cost experiment for comparison, we also finetune the model to transfer language capabilities. Fine-tuning an image generation model can often lead to catastrophic forgetting, which significantly degrade the model’s open-domain performance. To address these issues, we adopt a parameter-efficient fine-tuning strategy that preserves the original model’s ability to the greatest extent possible. During the learning stage, we only update the network parameters required for learning language character generation while freezing other parameters. We adopt two fine-tuning schemes here. The first on is Fine-tuning K and V matrices of the attention laters in the UNet. As proposed in [16], when updating the mapping from given text to image distribution, only updating $W_{K},W_{V}$ in each cross-attention block $i$ is sufficient since text features are the only input to the key and value projection matrix. The second one is fine-tuning with LoRA [12].
Evaluation only $W_{k}^{(i)}$ and $W_{v}^{(i)}$ LoRA PEA-Diffusion
CLIPScore_zh 0.2485 0.2366 0.2610
ImageReward 0.0293 -0.3214 1.0245
CLIPScore_en 0.3854 0.3577 0.4224
PickScore 0.1758 0.1613 0.2244
10 Performance of LLMs as Text Encoder for T2I
Table 8: Performance of LLMs as text encoder for T2I in Chinese specific language.
Model Base X-encoder ImageReward PickScore CLIPScore_en CLIPScore_zh
SDXL ChatGLM-6B 0.8547 0.2105 0.3985 0.2454
Qwen-7B 0.7547 0.2005 0.3918 0.2398
Baichuan2-7B 0.7952 0.2014 0.3907 0.2354
Palyground V2 ChatGLM-6B 0.8924 0.2154 0.4014 0.2415
Kandinsky V3 ChatGLM-6B 0.6751 0.1928 0.4032 0.2401
Stable Cascade ChatGLM-6B 0.6450 0.2014 0.3987 0.2584
To further enhance the performance of PEA-Diffusion on non-English text, we attempted to replace the encoder with a stronger X-Test encoder. Large Language Models (LLMs) have demonstrated exceptional semantic understanding and generation capabilities. Although LLMs are mostly designed as decoder-only structures, our experiments on Chinese indicate that LLMs can also serve as text encoders for SD through PEA adaptation. Table 8 presents the experimental results. We experimented with three different base models and four large language models.
Figure 13 showcases the effects of different models used as T2I text encoders, with the emphasize that the exhibited results are randomly selected. The first column is teacher model ProtoVision XL that based on SDXL and undergoes high-quality fine-tuning. The second column corresponds to PEA diffusion trained based on the Chinese CLIP. The subsequent three columns depict PEA diffusion trained with different LLMs. The eight prompts tested here are:
  • 3D Technological, Futuristic, Cyberpunk, Aurora, Smoke, Ice Dragon, Origami.
  • A cat with sunglasses holds a sword in his hand, in a demon castle, in the style of Sword and Fairy.
  • Grandiose cosmic backdrops, city ruins on floating islands, fantasy concept art, futurism.
  • Landscape, Clouds swirled around the house, Moon, World, Fantasy, mythology.
  • A winged woman in a white dress.
  • A steampunk cat with cool glasses, Portait, is all the rage in the art world, vaporwave art.
  • A very detailed, realistic photograph of a frog wearing small sunglasses and smoking a cigarette in a beautiful mid-century modern style chair, designed by Sonia Delaunay. Beautiful colors. Vintage film camera.
  • Godzilla became a professional boxer.
The four columns is the specific language input corresponding to X-Test encoder. Although there exists some disparity between the Imagereward in Table 8 and the teacher model, the qualitative subjective gap in visual effects is relatively small when viewing the overall presentation in Figure 13. The only exception is a few instances where the semantic alignment effects are slightly inferior, such as the last row of images in Figure 13.
Figure 13: More qualitative results with LLMs as text encoder for T2I.
Figure 14 illustrates the performance of different backbone frameworks when using ChatGLM-6B as the T2I encoder. The conclusions are similar to Figure 13 and further demonstrate the plug-and-play capabilities of the PEA method, allowing for rapid adaptation to different model frameworks.
Figure 14: Qualitative results of different model base. The implementation based on three different framework models demonstrates the flexible adaptability of PEA diffusion.
11 More Qualitative Results for Sec. 4.3 Table 1
Although the CLIPScore for MC in Sec. 4.3 Table 1 shows some improvement in specific language, the comparison of MG data against the performance of the teacher model is not particularly ideal. The ImageReward and PickScore metrics both show a slight decrease. Therefore, it is necessary to analyze the qualitative results.
Firstly, for the model base results in Figure 15 targeting SD1.5, it is evident that our method has a significant advantage over Altdiffusion and GlueGen. Compared to the teacher model, there is a decrease of 0.1442 in the ImageReward metric, indicating a slight decline in image quality. However, this situation improves significantly when SDXL is used as the model base.
In Figure 16, it can be observed that both LORA and $W_{K},W_{V}$ fine-tuning show a noticeable decrease in image quality. However, the last two columns, based on PEA strategies, especially the last column employing Multilingual-CLIP as PEA diffusion, exhibit a difficulty in qualitatively discerning the decline in image quality, despite the ImageReward metric showing a decrease of 0.1525. The same reasoning applies to the first two columns in Figure 13.
Here, we speculate that beyond a certain threshold of the ImageReward metric, possibly around 1 or greater, the differences in the metric may not necessarily translate into significant differences in actual outcomes. Of course, we will further validate this speculation through additional experiments.
Figure 15: Qualitative result comparisons in Sec. 4.3 Table 1 with SD1.5 model base.
Figure 16: Qualitative result comparisons in Sec. 4.3 Table 1 with SDXL model base.
12 User Study
Table 9: User studies.
visual quality semantic alignment success rate
Methods MC MG MC MG MC MG
Translate 2.2267 3.6126 1.2981 3.2042 0.2577 0.8178
LoRA 0.6536 1.6889 0.8727 0.9202 0.1639 0.1590
PEA-Diffusion 2.8423 3.6186 2.3539 3.0643 0.4978 0.8089
We generated 2000 images, based on a total of 400 prompts from MG and MC, with each prompt producing 5 images to determine the success rate. We then established specific scoring criteria and assigned the human evaluation to over ten experienced annotators. Their results were averaged to form the final result. The first two indicators were graded on a scale ranging from 0 to 4, with five levels in total. The higher the score, the better the performance. As shown in Table 9, our method exhibits a significant advantage over translation results and LoRA in three metrics for the MC dataset, particularly in semantic alignment and success rate. For the MG dataset, it is worth highlighting that there is no disparity between our method and translation regarding visual quality, while both semantic alignment and success rate are only marginally lower. This further substantiates that our approach’s efficacy does not diminish on general data. More details can be found in Appendix  11.
13 Downstream Experiment
13.1 LCM Quantitative Analysis
Latent Consistency Models [20] have achieved impressive performance in accelerating T2I generative tasks, producing high-quality images with minimal inference steps. “lcm-lora-sdx” is a universal acceleration module based on SDXL that allows to reduce the number of inference steps to only between 2 and 8 steps.
@@ 13.1 LCM Quantitative Analysis
13.2 More Downstream Generated Images
Figure 18 shows more image generation situations of downstream adaptation such as fine-tuned checkpoint, LoRA, ControlNet, Inpainting, SSD, LCM, SDXL-Turbo. All the pretrained checkpoints will be detailed in Appendix 6. Through Figure 18, we have proven that the PEA module can be used as a plugin between the Chinese CLIP text encoder and the UNet, achieving language transfer ability from English to Chinese without any other process.